OpenStudy (anonymous):

The lesson is positive exponents. This is the problem that needs to be simplified: (2^3)^2. In what occasion would I apply the outer exponent to the whole number?

OpenStudy (imstuck):



OpenStudy (theeric):

Try to remember the rules, but know there's hope if you forget! \[\left(2^3\right)^2\\ =\\ \left(2^3\right)\left(2^3\right)\\ =\\ \left(2\times2\times2\right)\left(2\times2\times2\right)\\ =\\ 2\times2\times2\times2\times2\times2\\ =\\ 2^6\]

OpenStudy (theeric):

Hopefully you just remember they can be multiplied. If not... Hopefully you can stop at \((2^3)^2=2^32^3\), because you know the right side exponents can be added. If not... Hopefully you can stop at \(2^32^3=2\times2\times2\times2\times2\times2\) because that is the exact reason why we have exponents! Counting how many times a factor multiplies itself!

OpenStudy (anonymous):

So in this case, do I only multiply the exponents?